tanx+tan(π+π/4)=2 find the most general value of x satisfying the given equation.

Grade:11

3 Answers

Aditya Gupta
5 years ago
tan(π+π/4)=1,
so we get tanx=2-1=1,
or tanx=tan(π/4),
or x=nπ+π/4, where n belongs to integer, which is the general solution
Arun
5 years ago
Dear Gulshan
 
tan x + tan pi/4 = 2
 
tan x = 2 – 1
 tan x = 1
 
hence
 
x = n * pi + pi/4
 
Regards
Arun (askIITians foryum expert)
Yash
5 years ago
Given,
TanX+Tan(π+π/4)=2
Tan( π+π/4)=1,
As we know that Tan(π+π/4)= 1
Then
TanX+1=2
 
TanX=2-1
 
         =1
TanX =1 it satisfy'' only when X  value is 45. Because Tan45=1